# Trends and Reversion in Financial Markets on Time Scales from Minutes to Decades

### v1

@@ 3.2 First Look at the Data
As for daily returns, we first visualize the data to decide how to set up the regression. We first focus on trend strengths in the range $[-2.5,+2.5]$ and the 3 short time horizons 2, 4, and 8 minutes. Fig. 3 (left) shows the scatter plots of the average next-minute return as a function of the trend strength at the end of the minute before, based on trend strength buckets as for daily data. The results are aggregated over all years and across all assets.
Figure 3: a: Average next-minute return as a function of current trend strength for $k=1,2,3$ (based on trend strength buckets). Left: the discontinuity at the origin due to the tick size is modelled by a step function. Right: the remaining pattern is modelled by a polynomial.
The average next-minute return is mostly positive for negative trend strength and negative for positive trend strength, so futures prices tend to revert at minute-scales. There is a discontinuity at trend strength 0. We have found that it does not go away if we make the trend strength buckets even smaller, and that it decreases with the trend horizon $T$ like $1/\sqrt{T}$. This discontinuity has a natural interpretation: futures prices are not continuous but discrete with a certain tick size. When the (compounded) underlying market price is in the middle between two ticks, the futures price tends to oscillate between them, leading to a negative auto-correlation of market returns from one minute to the next.
The size of the disontinuity depends on the tick size and the liquidity of the market. In the trend strength for horizon $T$, the weight of the previous minute is of order $1/\sqrt{T}$, which explains the decrease of the discontinuity with $T$. This effect can be used to reduce trading costs, but it can hardly be used for a profitable trading strategy. One could of course try to earn a ”market maker fee” by placing buy orders at the lower price just after an up-tick, and placing sell orders at the higher price just after a down-tick. However, there are trading costs. Moreover, there is the risk that the price does not revert but runs away due to an underlying trend or a buy/sell order by a large investor. We assume that markets are efficient in the sense that any trading profit just compensates for this risk.
Figure 4: Left: extension of fig. 3 (left) to larger and smaller trend strengths after removing the discontinuity. The hyperbolic tangent does not capture the tails well for $T\leq 4$. Right: same graph for $T=$1 hour to 1 day. The pattern is modelled well by a cubic polynomial.
@@
For $|t|<2$, the linear and cubic terms are sufficient. However, their signs are inverted compared with the results from daily data: Even without the discontinuity due to the tick size, we see that small trends tend to mean-revert (negative linear coefficient), while the reversion becomes weaker for stronger trends (positive cubic coefficient). A possible explanation is that small random fluctuations of the prices of the underlying are quickly corrected by statistical arbitrage traders. However, in cases where there is an underlying cause for the price change, such as an earnings announcement or an economic/political event, one expects that the trend becomes strong and preists until a new equilibrium price is reached that prices in this event. Such situations should weaken the general pattern of mean reversion at minute scales. Again, it would be difficult to profitably exploit these weak patterns in the face of trading costs, beyond earning a fair risk premium. In this sense, the patterns are still consistent with assuming that markets are efficient [22] at intraday scales.
@@ 3.3 Regression Analysis
We see from table 4 that the discontinuity $e$ at the origin due to the nonzero tick size is statistically highly significant. The linear coefficient $b$ is also highly significant, but now it is negative. This quantifies our qualitative observation from the scatter plot and confirms that, overall, markets tend to mean-revert at intraday scales even after discounting for the reversion due to the tick size. The cubic coefficient $c$ is also highly significant and positive, while it was highly significant (and negative) in the case of daily data. The quintic coefficient is negative and highly significant as well. Even terms were not highly signicficant.
Table 5 Intraday regression up to 16 minutes
@@
Next, we perform the regression aggregated across the longer time horizons $T=2^{k}$ minutes with $k\in\{6,7,8,9,10\}$, this time without the quintic term as discussed. The results are shown in table 6. We observe that the disontinuity has become smaller, as expected. The linear and cubic coefficients have switched sign and are now in line with the pattern from daily data: weak trends tend to persist ($b>0$), while strong trends tend to revert ($c<0$).
Table 6 Intraday regression $\geq 1$ hour
@@ 3.4 Refined Analysis by by Time Horizon
Figure 5: Regression coefficients for time scales of $T=2^{k}$ minutes, rescaled by $\sqrt{T}$. Left: discontinuity due to tick size and quintic term for short time horizons, rescaled for better visualization. Right: remaining linear and cubic coefficients. Dashed lines: same coefficients with a 1-minute time lag between trend measurement and predicted return.
@@
$\tilde{a}_{T}=a_{T}\cdot\sqrt{T\over 60}\ ,\ \tilde{b}_{T}=b_{T}\cdot\sqrt{T\over 60}\ ,\ \tilde{c}_{T}=c_{T}\cdot\sqrt{T\over 60}\ ,\ \tilde{d}_{T}=d_{T}\cdot\sqrt{T\over 60}\ ,\ \tilde{e}_{T}=e_{T}\cdot\sqrt{T\over 60}$
@@
$R_{T}(t+{T\over 60})\equiv\ln P(t+{T\over 60})-\ln P(t)\ \ \Rightarrow\ \ \langle R_{T}^{2}\rangle={T\over 60}\langle R^{2}\rangle=1.$
@@
We observe that the rescaled tick size discontinuity is indeed negative and almost independent of $T$, although it becomes statistically less significant for larger $k$ (fig. 5, left). The quintic term is statistically significant and negative for time scales less than one hour. Fig. 5 (right) shows that the regime where markets revert more strongly than implied by the tick size effect (negative $b$, positive $c$) ends at time scales of approximately 30 minutes. For longer time horizons, weak trends tend to persist (positive $b$, negative $c$). $b$ becomes larger than $|e|$ at scales of a few hours, indicating that these trends can even overcome the tick size reversion. This ”trending” regime extends up to daily scales, where our results connect with our previous results based on daily data [19], as seen in fig. 10 of section 5.
@@ 3.5 Refined Analysis by Asset Class and Time Zone
Fig. 6 shows the results for longer time scales $k=6,...,10$ (top) and for shorter time scales $k=1,2,3,4$ (bottom). The white marks represent the overall values of the regression parameters, followed by their values for (i)-(vii) in this order (black marks). Some regression parameters have been rescaled to fit all of them in the same range. We observe that for horizons of one hour or more (longer time scales), the different values of each parameter more or less agree within the estimation errors. In this sense, the parameter values may be universal. On the other hand, for time scales up to 16 minutes, the parameter values signifficantly differ from each other, depending on the asset class or time zone. Thus, at very short time scales, the parameter values are not universal. We return to this in scetion 5.
4 Century Time Scales
@@
4.1 Data
Table 7 1 2 3 4 5 6
@@ 4.1 Data
There are many issues with these data that have the potential to distort the results of our analysis. Let us illustrate them at a few examples.
@@
  • Apart from several equity market crashes, fig. 7 (left) shows a brief spike of the index around 1720, where its value briefly grew almost 10-fold, then crashed. This is not a data error. At that time the index consisted of only 3 companies, one of which - the South Sea Company - caused a huge speculative bubble that burst soonafter. Such outliers can also distort our analysis of shorter-term trends.
@@
In the subsequent analysis, we use trend horizons of $1.5\cdot 2^{k}$ months, with $k\in\{1,2,...,9\}$, corresponding to 3 months to 64 years. We use two data sets: one set of the monthly returns of the assets in table 8 since 1871, covering 150 years. These data are more reliable than older historical data, and include all of the markets in table 8 during the whole time period, except for the Nikkei index and for the shorter-term trends of currencies. We use them to analyze trends up to $k=7$ (16 years). For $k=7,8$ (32, 64 years), we include each market in table 8 since the start of its track record. The longest track record, that of the British stock market index, begins in 1692. These data are less reliable but contain just enough history to study trends that last a few decades.
@@ 4.2 First Look at the Data
To explore what kind of pattern we can expect, we again start with a visualization of the data based on trend strength buckets, as for daily and minute data. Fig. 8 (left) shows scatter plots of the average next-month return (on a log scale) as a function of a given month’s trend strength for shorter-term trends. We average over all years, all markets excluding currencies, and time horizons from 3 months to 1 year. We clearly see the cubic polynomial that we have already observed for daily data since 1990 over these trend horizons. The positive linear and negative cubic coefficients imply that weak trends tend to persist, while too strong trends tend to revert. The new piece of information we gain is that this cubic pattern is not a recent phenomenon, but has existed for at least 150 years, in line with previous results [16, 17].
@@
Figure 8: a: Average next-month return as a function of current trend strength (based on trend strength buckets). Left: horizons up to 2 years, Right: horizons of more than 2 years.
Fig. 8 (right) shows the same plot (including currencies) for long time hozions (4-16 years). The slope $b$ at the origin is clearly negative. The cubic coefficient $c$ seems to be positive. We will confirm this by the regression analysis of the next sub-section; we also find that the quadratic term is not significant. Thus, the scatter plot suggests that, for time horizons of many years, weak trends tend to revert. A possible interpretation is that this reflects economic cycles. On the other hand, strong long-term trends tend to persist. This may reflect the long-term rise and decline of individual economies, or of the supply and demand for key commodities such as oil or copper.
@@ 4.3 Regression Analysis
We now carry out two cubic regression analyses with monthly returns, one for shorter trend horizons of $T\in\{3,6,12\}$ months with all assets excluding currencies (to avoid the hyperinflation outliers), and one for longer horizons $T\in\{2,4,8,16\}$ years including currencies. We begin with the period 1871-1923.
Table 8 Monthly regression, up to 1 year Monthly regression, $2-16$ years
@@
$R^{2}$ $10.4$ bp 14.9 bp $4.5$ bp 7.8 bp
$R^{2}_{adj}$ $4.7$ bp 7.2 bp $1.5$ bp 3.3 bp
The results are shown in table 8. They confirm and quantify our qualitative observations that the linear coefficient $b$ is positive for shorter horizons and negative for longer horizons, while the cubic coefficient $c$ is negative for shorter horizons and positive for longer horizons. These results are statistically significant; the errors are computed by bootstrapping with 200 samples. The adjusted R-squared is computed by 12-fold cross validation. We have also verified that the quadratic term is not significant for shorter time horizons. For longer horizons, it is weakly significant, but we leave it out to keep the model more robust.
Figure 9: Regression of monthly returns against trend strengths by time horizon. Left: all assets since 1871, excluding currencies for $T\leq 1$ year. Right: extension of the analysis to 6 decades, based on each asset since its track record starts ($\geq 1692$).
@@ 4.4 Refinements and Extension to Medieval Times
To check if our reversion pattern for long trend horizons depends on the asset class, we have performed seperate nonlinear regression analyses for equity indices, interest rates, commodities, and currencies. As we have already found in the case of daily data, for time horizons up to 1 year all asset classes show the same pattern: weak trends tend to persist, strong trends tend to revert. Unfortunately, for longer-term horizons, the amount of noise in our monthly long-term data does not allow us to establish whether the pattern found above (reversion of weak trends, persistence of strong trends) occurs seperately for each asset class.
However, on qualitative grounds, it is clear that there are differences between the asset classes. Equities differ by their long-term risk premia of around 5% per annum. Interest rates have generally declined over the centuries and decades. Currencies have extreme down-trends in times of hyper-inflation, and commodity prices display periods of extreme volatility. Thus, the universality of the interplay of trends and reversion that we have observed from daily data since 1990 may not extend to time horizons of a few years and longer.
It would also be interesting to see how the interplay of trends and reversion has evolved over the centuries, extending linear results in [16, 17, 18]. Unfortunately, our data set is too limited to yield statistically significant nonlinear regression results for earlier centuries.
We have also tried to extend our observations on the interplay of trends and reversion all the way back to medieval times. We could do this only for interest rates, for which we have annual quotes for Germany (since 1327), France (since 1388), England (since 1311), and the Netherlands (since 1401). There are also historical prices for gold and silver since 1258, but since gold and silver essentially were the currencies at the time, these prices are not useful for a regression analysis: they tend to be stable over decades, interrupted by occasional jumps in times of inflation. Price quotes for more exotic assets such as silk or safran are too infrequent and too discontinuous to be included in the analysis.
Table 9 Yearly linear regression for interest rates since the 14th century
@@
A simple linear regression for interest rates since the 14th century, with trend horizons ranging from 2 to 128 years, does confirm, not surprisingly, that trends in interest rates that last several decades tend to revert, as shown in table 9. However, the limited amount of such data that go back to medieval times does not permit a more refined nonlinear regression.

### v2

@@ 3.2 First Look at the Data
As for daily returns, we first visualize the data to decide how to set up the regression. We first focus on trend strengths in the range $[-2.5,+2.5]$ and on the 3 short time horizons 2, 4, and 8 minutes. Fig. 3 (left) shows the scatter plots of the average next-minute return as a function of the trend strength at the end of the minute before, based on trend strength buckets as for daily data. The results are aggregated over all years and across all assets.
Figure 3: Average next-minute return as a function of the current trend strength for horizon $T=2^{k}$ minutes with $k=1,2,3$. Left: the discontinuity at the origin due to the tick size is modelled by a step function. Right: the remaining pattern is modelled by a polynomial.
The average next-minute return is mostly positive for negative trend strength and negative for positive trend strength, so futures prices tend to revert at minute-scales. There is a discontinuity at trend strength 0. We have found that it does not go away if we make the trend strength buckets even smaller, and that it decreases with the trend horizon $T$ like $1/\sqrt{T}$. This discontinuity has a natural interpretation: futures prices are not continuous but discrete with a certain tick size. E.g., S&P 500 futures are priced in units of 25 cent (e.g., at \$5012.00 or \$5012.25, but not in between). When the (compounded) underlying market price is in the middle between two ticks, the futures price tends to oscillate between them, leading to a negative auto-correlation of market returns from one minute to the next.
The size of the discontinuity depends on the tick size and the liquidity of the market. In the trend strength for horizon $T$, the weight of the previous minute is of order $1/\sqrt{T}$, which explains the decrease of the discontinuity with $T$. This effect can be used to reduce trading costs, but it can hardly be used for a profitable trading strategy. One could of course try to earn a so-called ”market maker fee” by placing buy orders at the lower price just after an up-tick, and placing sell orders at the higher price just after a down-tick. However, there are trading costs. Moreover, there is the risk that the price does not revert but runs away due to an underlying trend or a buy/sell order by a large investor. We assume that markets are efficient in the sense that any trading profit just compensates for this risk.
Figure 4: Left: extension of fig. 3 (left) to larger and smaller trend strengths after removing the discontinuity. A hyperbolic tangent models the tails poorly for horizons $T\leq 4$. Right: analogous graph for $T=1$ hour to 1 day. The pattern is modelled well by a cubic polynomial.
@@
For $|t|<2$, the linear and cubic terms are sufficient. However, their signs are inverted compared with the results from daily data: Even without the discontinuity due to the tick size, we see that small trends tend to mean-revert (negative linear coefficient), while the reversion becomes weaker for stronger trends (positive cubic coefficient).
A possible explanation is that small random fluctuations of the prices of the underlying are quickly corrected by statistical arbitrage traders (i.e., traders who systematically sell/buy assets whose price has risen/fallen significantly, thereby driving the asset price back down/up). However, in cases where there is an underlying cause for the price change, such as an earnings announcement or an economic/political event, one expects that the trend becomes strong and persists until a new equilibrium price is reached that prices in this event. Such situations should weaken the general pattern of mean reversion at minute scales. Again, it would be difficult to profitably exploit these weak patterns in the face of trading costs, beyond earning a fair risk premium. In this sense, the patterns are still consistent with assuming that markets are efficient [26] at intraday scales.
@@ 3.3 Regression Analysis
Parameter Regression results for trend horizons up to 16 minutes
@@
Table 5: Regression results with minute data according to ansatz (9) for short-term trends.
Parameter Regression results for trend horizons $\geq 1$ hour.
@@
Table 6: Regression results with minute data according to ansatz (9) for longer-term trends.
We see from table 5 that the discontinuity $e$ at the origin due to the nonzero tick size is statistically highly significant. The linear coefficient $b$ is also highly significant, but now it is negative. This quantifies our qualitative observation from the scatter plot and confirms that, overall, markets tend to mean-revert at intraday scales even after discounting for the reversion due to the tick size. The cubic coefficient $c$ is also highly significant and positive, while it was negative in the case of daily data. The quintic coefficient is negative and highly significant as well. Even terms were not significant enough to be included in the analysis, as explained in the appendix.
Next, we perform the regression aggregated across the longer time horizons $T=2^{k}$ minutes with $k\in\{6,7,8,9,10\}$, this time without the quintic term as discussed. The results are shown in table 6. We observe that the discontinuity has become smaller, as expected. The linear/cubic coefficients have switched signs and are now in line with the pattern from daily data: weak trends tend to persist ($b>0$), while strong trends tend to revert ($c<0$).
@@ 3.4 Refined Analysis by by Time Horizon
$(\tilde{a}_{T},\tilde{b}_{T},\tilde{c}_{T},\tilde{d}_{T},\tilde{e}_{T})\ =\ (a_{T},b_{T},c_{T},d_{T},e_{T})\cdot\sqrt{T\over 60}$ (11)
@@
$R_{T}(t+{T\over 60})\equiv\ln P(t+{T\over 60})-\ln P(t)\ \ \Rightarrow\ \ \langle R_{T}^{2}\rangle={T\over 60}\langle R^{2}\rangle=1.$ (12)
@@
Figure 5: Regression coefficients in ansatz (10) for trend horizons $T=2^{k}$ minutes, rescaled by $\sqrt{T}$. Left: discontinuity $\tilde{e}_{T}$ due to tick size and quintic coefficient $\tilde{d}_{T}$ for short time horizons (rescaled by factors). Right: remaining linear and cubic coefficients $\tilde{b}_{T}$ and $\tilde{c}_{T}$. Dashed lines: same coefficients with a 1-minute time lag between measured trend and predicted return.
We observe in fig. 5 that the rescaled tick size discontinuity is indeed negative and almost independent of $T$, although it becomes statistically less significant for larger $k$ (fig. 5, left). The quintic term is statistically significant and negative for time scales less than one hour. Fig. 5 (right) shows that the regime where markets revert more strongly than implied by the tick size effect (negative $b$, positive $c$) ends at time scales of approximately 30 minutes. For longer time horizons, weak trends tend to persist (positive $b$, negative $c$). $b$ becomes larger than $|e|$ at scales of a few hours, indicating that these trends can even overcome the tick size reversion. This ”trending” regime extends up to daily scales, where our results connect with our previous results based on daily data [23], as seen in fig. 10 of section 5.
@@ 3.5 Refined Analysis by Asset Class and Time Zone
Fig. 6 shows the results for longer time scales $k=6,...,10$ (top) and for shorter time scales $k=1,2,3,4$ (bottom). The white marks represent the overall values of the regression parameters, followed by their values for (i)-(vii) in this order (black marks). Some regression parameters have been rescaled to fit all of them in the same range. We observe that for horizons of one hour or more (longer time scales), the different values of each parameter more or less agree within the estimation errors. In this sense, the parameter values may be universal. On the other hand, for time scales up to 16 minutes, the parameter values significantly differ from each other, depending on the asset class or time zone. Thus, at very short time scales, the parameter values are not universal. We return to this in scetion 5.
4 Analysis of Centuries of Data
@@
Trends in past centuries have previously been analyzed. In [12], it was found that trend-following on time scales from 1 month to 1 year would have been highly profitable across all asset classes since the 1880. In [13], this result was extended back to 1800. According to [14], trends may even have existed for at least 800 years. After the analysis in this section was completed, we noticed that [15] extends the 200-year analysis of [13] to trend horizons of up to 5 years, including the cubic and also a quadratic term in ansatz (1). The independently obtained results that we report below are consistent with those of [15]. They extend them to more than 3 centuries, and to trend horizons of several decades.
4.1 Data Set
Asset Class 1 2 3 4 5 6
@@ 4.1 Data Set
Table 7: List of financial markets used in the analysis of monthly data covering centuries.
@@
There are many issues with these data that have the potential to distort the results of our analysis. Let us illustrate them at a few examples; see also the review of the same data set in [13].
@@
  • Apart from several equity market crashes, fig. 7 (left) shows a brief spike of the index around 1720, where its value briefly grew almost 10-fold, then crashed. This is not a data error. At that time the index consisted of only 3 companies, one of which - the South Sea Company - caused a huge speculative bubble that burst soonafter. Such extreme outliers can also distort our analysis of shorter-term trends.
@@
In view of these data issues with century-old data, in this section we will regard regression coefficients as statistically significant only if their t-statistics is $\geq 3$. We use trend horizons of $1.5\cdot 2^{k}$ months, with $k\in\{1,2,...,9\}$, corresponding to 3 months to 64 years. We use two data sets: one set of the monthly returns of the assets in table 8 since 1871, covering 150 years. These data are more reliable than older historical data, and include all of the markets in table 8 during the whole time period, except for the Nikkei index and for the shorter-term trends of currencies. We use them to analyze trends up to $k=7$ (16 years). For $k=7,8$ (32, 64 years), we include each market in table 8 since the start of its track record. The longest track record, that of the British stock market index, begins in 1692. These data are less reliable but contain just enough history to study trends that last a few decades.
@@ 4.2 First Look at the Data
Figure 8: Average next-month return as a function of the current trend strength, based on 150 years of data. Left: trend horizons up to 2 years. Right: horizons of more than 2 years.
To explore what kind of pattern we can expect, we again start with a visualization of the data by grouping them into buckets of similar trend strengths, as for daily and minute data. Here we use 5 buckets per unit of trend strength. Fig. 8 (left) shows scatter plots of the average next-month return (on a log scale) as a function of a given month’s trend strength for shorter-term trends. We average over all years, all markets excluding currencies, and time horizons from 3 months to 1 year.
We clearly see the cubic polynomial that we have already observed for daily data since 1990 over these trend horizons. The positive linear and negative cubic coefficients imply that weak trends tend to persist, while too strong trends tend to revert. The new piece of information we gain is that this cubic pattern is not a recent phenomenon, but has existed for at least 150 years, in line with previous results [12, 13, 15].
@@
Fig. 8 (right) shows the same plot (including currencies) for long time horizons (4-16 years). The slope $b$ at the origin is clearly negative. The cubic coefficient $c$ seems to be positive. We will confirm this by the regression analysis of the next sub-section. Thus, the scatter plot suggests that, for time horizons of many years, weak trends tend to revert. A possible interpretation is that this reflects economic cycles. On the other hand, strong long-term trends tend to persist. This may reflect the long-term rise and decline of individual economies, or of the supply and demand for key commodities such as oil or copper.
@@ 4.3 Regression Analysis
We now carry out two cubic regression analyses with ansatz (1) for monthly returns, one for shorter trend horizons of $T\in\{3,6,12\}$ months with all assets excluding currencies (to avoid the hyperinflation outliers), and one for longer horizons $T\in\{2,4,8,16\}$ years including currencies. We begin with the period 1871-1923.
Parameter Trend horizons up to 1 year Trend horizons $2-16$ years
@@
$R^{2}$ $11.0$ bp 15.4 bp $5.0$ bp 8.5 bp
$R^{2}_{adj}$ $5.7$ bp 8.4 bp $2.2$ bp 4.4 bp
Table 8: Regression results for ansatz (1), using long-term monthly (instead of daily) data. Left: shorter horizons excluding currencies. Right: longer horizons including currencies.
The results are shown in table 8. They confirm and quantify our qualitative observations that the linear coefficient $b$ is positive for shorter horizons and negative for longer horizons, while the cubic coefficient $c$ is negative for shorter horizons and positive for longer horizons. These results are statistically significant; the errors are computed by bootstrapping with 500 samples. The adjusted R-squared is computed by 12-fold cross validation. As reported in the appendix, a quadratic term would clearly be insignificant for shorter time horizons. For longer horizons, its $t$-statistics is $t=2.4$, which is below our significance threshold of 3. In view of this, we do not add a quadratic term, although we cannot rule out its existence.
Figure 9: Regression of monthly returns against trend strengths by time horizon, using ansatz (1). Left: all assets since 1871, excluding currencies for $T\leq 1$ year. Right: extension of the analysis to 6 decades, based on each asset since the first year of its track record ($\geq 1692$).
@@
Our results are in close agreement with the 200-year-results [15], where it was also found that the parameters we call $b_{T}$ and $c_{T}$ both switch signs at time horizon $T\approx 2$ years (although we disregard the negative quadratic term). Our results suggest that this reversion behavior extends from time scales of 5 years as in [15] to at least a few decades, and that it has existed for at least 300 years.
@@ 4.4 Refinements and Extension to Medieval Times
To check if our reversion pattern for long trend horizons depends on the asset class, we have performed separate nonlinear regression analyses for equity indices, interest rates, commodities, and currencies. As we have already found in the case of daily data, for time horizons up to 1 year all asset classes show the same pattern: weak trends tend to persist, strong trends tend to revert. Unfortunately, for longer-term horizons, the amount of noise in our monthly long-term data does not allow us to establish whether the pattern found above (reversion of weak trends, persistence of strong trends) occurs separately for each asset class.
However, on fundamental grounds, it is clear that there are differences between the asset classes. Equities differ by their long-term risk premia of around 5% per annum. Interest rates have generally declined over the centuries and decades. Currencies have extreme down-trends in times of hyper-inflation, and commodity prices display periods of extreme volatility. Thus, the universality of the interplay of trends and reversion that we have observed from daily data since 1990 is unlikely to extend to time horizons of a few years and longer.
We have also tried to extend our observations on the interplay of trends and reversion all the way back to medieval times, trying to extend linear results in [14], but our data set is too limited to yield statistically significant nonlinear regression results.
Parameter Regression results ($a=$ intercept, $b=$ beta to trend strength)
@@
Table 9: Linear regression results for yearly interest rates since the 14th century.
We could only do a simple linear analysis for interest rates, for which we have annual quotes for Germany (since 1327), France (since 1388), England (since 1311), and the Netherlands (since 1401). There are also historical prices for gold and silver since 1258, but since gold and silver essentially were the currencies at the time, these prices are not useful for a regression analysis: they tend to be stable over decades, interrupted by occasional jumps in times of inflation. Price quotes for more exotic assets such as silk or safran are too infrequent and too discontinuous to be included in the analysis. The linear regression for interest rates since the 14th century, with trend horizons from 2 to 128 years, is shown in table 9. Not surprisingly, it indicates that trends in interest rates that last several decades tend to revert.